For values of L = 1H; C=1nF; R = 5kΩ, what is the resonant frequency of the
circuit? What is the natural frequency of the circuit? What is the Q value? If the coil is
wound from 42 gauge copper wire, what is the length of wire used in the coil?
Fig Q1
Answer #1

Pupedance (7) = (R+52) // %ec Z - (R+SL)• Yọc 1. Risit the RIBL (Z) = - PRETS2LCHI = (Rj6L) CR-W2C) - jWRC) (1-w2Rc)2+(WRC)”.h 42 guage copper turns =) N = 1000 Diameter =) 0.0633 mm. 14D10 &12 =2f ( W = 31692.7766) / RT L = 1480.0639 X 1000 x 10?m 2